a water tank has an inlet pipe and a drain pipe. A full tank can be emptied in 30 minutes if the drain is opened and an empty tank can be filled in 45 minutes with the inlet pipe opened. If both pipes are accidentally opened when the tank is full, then how long will it take to empty the tank
Answer by edjones(7569)

flow*time=volume f*t=v f=v/t
let v=100 gallons actually it can be any number you choose and the answer will be the same.
100/30=10/3
100/45=20/9
10/3-20/9=90/27-60/27=30/27=10/9 (10 gal leaves the tank every 9 minutes) So in 90 min the tank will be empty.
if you like formulas you can do a proportion:
10/9=100/t
10x/9=100
t=100*9/10=90 min
Ed